OpenStudy (anonymous):

( ( x−1 ) ( y4 ) ) -------------- 3x−5 y−1 please help solve i keep getting 3x^4y^5 but i think the answer is x^4y^5 ------ 3

OpenStudy (ranga):

\[\Large \frac{a^m}{a^n} = a^{m-n}\]

OpenStudy (ranga):

\[\Large \frac{ x^{-1}y^4 }{ 3x^{-5}y^{-1} } = \frac{ x^{-1 + 5}y^{4+1} }{ 3 } = ?\]